OpenStudy (anonymous):

I really need help with all of those questions: 1) 3√18 + 5√50 2) 3√8 - √32 + 3√72 - √75 3)1/√7 4)√2/√5 I am having a hard time with this lesson can you please explain how to do each one of those? thank you

OpenStudy (anonymous):

yeah but i have to wrote how i got the answer

OpenStudy (anonymous):

\[3\sqrt{18}\] is the same as three times the square root of 18, treat each part containing a radical the same way.

OpenStudy (anonymous):

18 does not have a square root, so we have to break it down until we find a number that has a nice looking square root. 9 and 2 multiply to make 18, and the square root of 9 is three, so we can write the first part of problem one as \[3\sqrt{2}\], because the square root of nine is three, so it can come out, and the two left over remains under the radical.

OpenStudy (mimi_x3):

Um , i think that it can't be \[3\sqrt{2}\] its \[9\sqrt{2}\]

OpenStudy (anonymous):

oh thnx

OpenStudy (anonymous):

You sure, Mimi? and for the second part of that problem, fifty breaks down into 2 * 25, and the square root of 25 is 5, so the five comes out, and it looks like \[5\sqrt{2}\]. Notice how the numbers under the radicals are the same? These are called like bases, so now all you have to do is add the 3 and the 5, and your final answer should be \[8\sqrt{2}\] for problem one

OpenStudy (anonymous):

ok i got that what about number 2?

OpenStudy (anonymous):

same deal, break it down and do the arithmetic.

OpenStudy (anonymous):

eh its kind of more confusing

OpenStudy (anonymous):

I can give you a tip, the eight under the radical in part one looks like a solid base for a like term, so see what factors of eight you can find to square root and pull out.

OpenStudy (mimi_x3):

Yup im sure , super because when you \[3\sqrt{2}\] in the calculator its 4.24..... that doesn't equal to \[3\sqrt{18}\] but when you put \[9\sqrt{2}\] it equals to 12.727... that is the same as \[3\sqrt{18}\] Normally , wouldn't be sure but this time I am xD

OpenStudy (anonymous):

I have to disagree, nine can not be the square root of nine, it appears you are just shifting it out from under the radical.

OpenStudy (anonymous):

However, I did just catch an error I made, Multiply the three you pulled out with the three already there, and our answers match.

OpenStudy (mimi_x3):

Well, idk what radical means but try it in your calculator and you will see . but then the 9 is outside the square root so you don't sqaure root it again. So i was right , right ?

OpenStudy (anonymous):

right, but remember, you pulled out a three from under the radical, and then multiplied it to the three outside. A nine did not come from the radical, because the answer would be 27 times the square root of two. A radical is the "house" the 18 lives in.

OpenStudy (mimi_x3):

LOL, the 'house' ok , i learnt a new thing today xD , isn't that a square root then ?

OpenStudy (anonymous):

oh ok thnx ^_^